# Regret Minimization in Stackelberg Games with Side Information

### v2

@@ Abstract
In its most basic form, a Stackelberg game is a two-player game in which a leader commits to a (mixed) strategy, and a follower best-responds. Stackelberg games are perhaps one of the biggest success stories of algorithmic game theory over the last decade, as algorithms for playing in Stackelberg games have been deployed in many real-world domains including airport security, anti-poaching efforts, and cyber-crime prevention. However, these algorithms often fail to take into consideration the additional information available to each player (e.g. traffic patterns, weather conditions, network congestion), a salient feature of reality which may significantly affect both players’ optimal strategies. We formalize such settings as Stackelberg games with side information, in which both players observe an external context before playing. The leader then commits to a (possibly context-dependent) strategy, and the follower best-responds to both the leader’s strategy and the context. We focus on the online setting in which a sequence of followers arrive over time, and the context may change from round-to-round. In sharp contrast to the non-contextual version, we show that it is impossible for the leader to achieve good performance (measured by regret) in the full adversarial setting (i.e., when both the context and the follower are chosen by an adversary). However, it turns out that a little bit of randomness goes a long way. Motivated by our impossibility result, we show that no-regret learning is possible in two natural relaxations: the setting in which the sequence of followers is chosen stochastically and the sequence of contexts is adversarial, and the setting in which the sequence of contexts is stochastic and the sequence of followers is chosen by an adversary.
@@ 1 Introduction
A Stackelberg game [32, 33] is a strategic interaction between two utility-maximizing players in which one player (the leader) is able to commit to a (possibly mixed) strategy before the other player (the follower) takes an action. While Stackelberg’s original formulation was used to model economic competition between firms, Stackelberg games have been used to study a wide range of topics in computing ranging from incentives in algorithmic decision-making [17] to radio spectrum utilization [34]. Perhaps the most successful application of Stackelberg games to solve real-world problems is in the field of security, where the analysis of Stackelberg security games has led to new methods in domains such as passenger screening at airports [9], wildlife protection efforts in conservation areas [12], the deployment of US Federal Air Marshals on board commercial flights [20], and patrol boat schedules for the US Coast Guard [2]. [^1] See [29, 21, 3] for an overview of other application domains for Stackelberg security games.
@@ 1.1 Overview of Our Results
We show that in the fully adversarial setting (i.e. the setting in which both the sequence of contexts and follower types is chosen by an adversary), no-regret learning is not possible. We show this via a reduction from the problem of online linear thresholding, for which it is known that no no-regret learning algorithm exists. Motivated by this impossibility result, we study two natural relaxations: (1) a setting in which the sequence of contexts is chosen by an adversary and the sequence of follower types is chosen stochastically (the stochastic follower setting), and (2) a setting in which the sequence of contexts is chosen stochastically and the sequence of follower types is chosen by an adversary (the stochastic context setting).
@@
In the stochastic follower setting we show that the greedy algorithm (Algorithm 1), which estimates the leader’s expected utility for the given context and plays the mixed strategy which maximizes their estimate, achieves no-regret as long as the total variation distance between their estimate and the true distribution is decreasing with time. We then show how to instantiate the leader’s estimation procedure so that the regret of Algorithm 1 is $O(\min\{K,A_{f}\}\sqrt{T\log(T)})$, where $T$ is the time horizon, $K$ is the number of follower types, and $A_{f}$ is the number of follower actions. In the stochastic context setting, we show the leader can obtain $O(\sqrt{KT\log(T)}+K)$ regret by playing Hedge over a finite set of policies (Algorithm 2). An important intermediate result in both settings is that it is (nearly) without loss of generality to consider leader policies which map to a finite set of mixed strategies $\mathcal{E}_{\mathbf{z}}$, given context $\mathbf{z}$ [^2] Specifically, we show that a leader who is restricted to playing such policies incurs negligible regret when compared to a leader who is not.
@@
Next, we extend our algorithms for both types of adversary to the setting in which the leader does not get to observe the follower’s type after each round, but instead only gets to observe their action. We refer to this type of feedback as bandit feedback. Both of our extensions to bandit feedback make use of the notion of a barycentric spanner [5], a special basis under which bounded loss estimators may be obtained for all leader mixed strategies. In the bandit stochastic follower setting, we use the fact that in addition to being bounded, a loss estimator constructed using a barycentric spanner has low variance, in order to show that a natural extension of our greedy algorithm obtains $O(T^{2/3})$ regret. [^3] We suppress $A_{f}$, $K$, and logarithmic terms for brevity. We also make use of barycentric spanners in the (bandit) stochastic context setting, albeit in a different way. Here, our algorithm proceeds by splitting the time horizon into blocks, and using a barycentric spanner to estimate the leader’s utility from playing according to a set of special policies in each block. We then play Hedge over these policies to obtain $O(T^{2/3})$ leader regret. [^4] This is similar to how Balcan et al. [6] use barycentric spanners to obtain no-regret in the non-contextual setting, although more care must be taken to handle the side information present in our setting. We discuss similarities to their work in more detail in Section 1.2. See Table 1 for a summary of our results.
@@
Stochastic Followers, Adversarial Contexts $O\left(\min\{K,A_{f}\}\sqrt{T\log T}\right)$ (Section 4.1) $O\left(K^{2/3}A_{f}^{2/3}T^{2/3}\log^{1/3}T\right)$ (Section 5.1)
Stochastic Contexts, Adversarial Followers $O\left(\sqrt{KT\log T}+K\right)$ (Section 4.2) $O\left(KA_{f}^{1/3}T^{2/3}\log^{1/3}T\right)$ (Section 5.2)
@@ 1.2 Further Related Work
Conitzer and Sandholm [10] provide algorithms for recovering the leader’s optimal mixed strategy in Stackelberg games when the follower’s payoff matrix is known to the leader. Letchford et al. [23] consider the problem of learning the leader’s optimal mixed strategy in the repeated Stackelberg game setting against a perfectly rational follower with an unknown payoff matrix. Peng et al. [25] study the same setting as Letchford et al. [23]. They provide improved rates and prove nearly-matching lower bounds. Learning algorithms to recover the leader’s optimal mixed strategy have also been studied in Stackelberg security games [6, 8, 28].
Our work builds off of several results established for online learning in (non-contextual) Stackelberg games in Balcan et al. [6]. In particular, our results in Section 4.2 and Section 5.2 may be viewed as a generalization of their results to the setting in which the payoffs of both players depend on an external context, in addition to the player actions. Roughly speaking, Balcan et al. [6] show that it without loss to play Hedge over a finite set of mixed strategies in order to obtain no-regret against an adversarially-chosen sequence of follower types. In order to handle the additional side information available in our setting, we instead play Hedge over a finite set of policies, each of which map to a finite set of (context-dependent) mixed strategies. However, the discretization argument is more nuanced in our setting. In particular, it is not without loss of generality to consider a finite set of policies. As a result, we need to bound the additional regret incurred by the leader due to the discretization. More recent work on learning in Stackelberg games provides improved regret rates in the full feedback [11] and bandit feedback [7] settings, and considers the effects of non-myopic followers [15] and followers who respond to calibrated predictions [16].
@@ 2 Setting and Background
We use $[N]:=\{1,\ldots,N\}$ to denote the set of natural numbers up to and including $N\in\mathbb{N}$ and cl$(\mathcal{P})$ to denote the closure of the set $\mathcal{P}$. $\mathbf{x}[a]$ denotes the $a$-th component of vector $\mathbf{x}$, and $\Delta(\mathcal{A})$ denotes the probability simplex over the set $\mathcal{A}$.
@@
is the total variation distance between distributions $\mathbf{p}$ and $\mathbf{q}$, and $\mathbb{E}_{t}[x]=\mathbb{E}[x|\mathcal{F}_{t}]$ is shorthand for the expected value of the random variable $x$, conditioned on the filtration up to (but not including) time $t$. All proofs which are not in the main body may be found in the Appendix. Finally, while we present our results for general Stackelberg games with side information, our results are readily applicable to the special case of Stackelberg security games with side information.
@@
We consider a repeated Stackelberg game between a leader and a sequence of followers. In each round $t\in[T]$, Nature reveals a context $\mathbf{z}_{t}\in\mathcal{Z}\subseteq\mathbb{R}^{d}$ to both players. [^5] For example, in airport security settings, $\mathbf{z}_{t}$ could be a hand-crafted feature vector containing information about arrival and departure times, number of passengers, valuable cargo, etc. In cyber-defense settings, $\mathbf{z}_{t}$ may be a list of network traffic statistics. The leader moves first by playing some mixed strategy $\mathbf{x}_{t}\in\mathcal{X}\subseteq\mathbb{R}^{A}$ over a set of (finite) leader actions $\mathcal{A}$, i.e., $\mathbf{x}_{t}\in\Delta(\mathcal{A})$. We denote the size of $\mathcal{A}$ by $A:=|\mathcal{A}|$. Having observed the leader’s mixed strategy, the follower best-responds to both $\mathbf{x}_{t}$ and $\mathbf{z}_{t}$ by playing some action $a_{f}\in\mathcal{A}_{f}$, where $\mathcal{A}_{f}$ is the (finite) set of follower actions and $A_{f}:=|\mathcal{A}_{f}|$. [^6] It is without loss of generality that the follower’s best-response is a pure strategy.
@@
Follower $f$’s best-response to context $\mathbf{z}$ and mixed strategy $\mathbf{x}$ is
$b_{f}(\mathbf{z},\mathbf{x})\in\arg\max_{a_{f}\in\mathcal{A}_{f}}\sum_{a_{l}\in\mathcal{A}}\mathbf{x}[a_{l}]\cdot u_{f}(\mathbf{z},a_{l},a_{f}),$
where $u_{f}:\mathcal{Z}\times\mathcal{A}\times\mathcal{A}_{f}\rightarrow[0,1]$ is follower $f$’s utility function. In the case of ties, we assume that there is a fixed and known ordering over actions which determines how the follower best-responds, i.e. if $a>a^{\prime}$ for $a,a^{\prime}\in\mathcal{A}_{f}$ then the follower will break ties between $a$ and $a^{\prime}$ in favor of $a$.
We allow for the follower in round $t$ (denoted by $f_{t}$) to be one of $K\geq 1$ follower types $\{\alpha^{(1)},\ldots,\alpha^{(K)}\}$ (where $K\leq T$). Follower type $\alpha^{(i)}$ is characterized by utility function $u_{\alpha^{(i)}}:\mathcal{Z}\times\mathcal{A}\times\mathcal{A}_{f}\rightarrow[0,1]$, i.e. given a context $\mathbf{z}$, leader action $a_{l}$, and follower action $a_{f}$, a follower of type $\alpha^{(i)}$ would receive utility $u_{\alpha^{(i)}}(\mathbf{z},a_{l},a_{f})$. We assume that the set of all possible follower types and their utility functions are known to the leader, but that the follower’s type at round $t$ is not revealed to the leader until after the round is over. We denote the leader’s utility function by $u:\mathcal{Z}\times\mathcal{A}\times\mathcal{A}_{f}\rightarrow[0,1]$ and assume it is known to the leader. We often use the shorthand
$u(\mathbf{z},\mathbf{x},a_{f})=\sum_{a_{l}\in\mathcal{A}}\mathbf{x}[a_{l}]\cdot u(\mathbf{z},a_{l},a_{f})$
to denote the leader’s expected utility of playing mixed strategy $\mathbf{x}$ under context $\mathbf{z}$ against follower action $a_{f}$. Follower $t$’s expected utility $u_{f_{t}}(\mathbf{z},\mathbf{x},a_{f})$ is defined analogously.
A leader policy $\pi:\mathcal{Z}\rightarrow\mathcal{X}$ is a (possibly random) mapping from contexts to mixed strategies. If the leader using policy $\pi_{t}$ in round $t$ and observes context $\mathbf{z}_{t}$, their mixed strategy $\mathbf{x}_{t}$ is given by $\mathbf{x}_{t}:=\pi_{t}(\mathbf{z}_{t})$.
@@
If an algorithm achieves regret $R(T)=o(T)$ (i.e. regret grows sublinearly as a function of $T$), we say that it is a no-regret algorithm.
In the sequel, we consider three ways in which Nature can select the sequence of contexts and/or followers:
@@
    If Nature chooses context $\mathbf{z}_{t}$ (resp. follower $f_{t}$) before round $t$ in order to harm the leader (possibly using knowledge of the leader’s algorithm and the outcomes of the prior $t-1$ roundes), we say that the sequence of contexts (resp. follower types) are chosen by an adaptive adversary.
@@ 3 On the Impossibility of Fully Adversarial No-Regret Learning
We begin with a negative result: no-regret learning is not possible in the setting of Section 2 if the sequence of contexts and the sequence of followers is chosen by an adversary. We show this result via a reduction to the online linear thresholding problem, for which it is known that no algorithm can achieve no-regret. In particular, we show that if there exists a no-regret algorithm for the leader in the fully adversarial version of our setting, then it could be used to construct a no-regret algorithm for the online linear thresholding problem.
Background on Online Linear Thresholding.
The online linear thresholding problem is a repeated two-player game between a learner and an adversary. Before the first round, an adversary chooses a cutoff $s\in[0,1]$ which is unknown to the learner. In each round, the adversary chooses a point $\omega_{t}\in[0,1]$ and reveals it to the learner. $\omega_{t}$ is assigned label $y_{t}=1$ if $\omega_{t}>s$ and label $y_{t}=-1$ otherwise. Given $\omega_{t}$, the learner makes a guess $g_{t}\in[0,1]$ (the probability they place on $y_{t}=1$), and receives utility
$u_{\mathrm{OLT}}(\omega_{t},g_{t})=g_{t}\cdot\mathbbm{1}\{y_{t}=1\}+(1-g_{t})\cdot\mathbbm{1}\{y_{t}=-1\}.$
The learner gets to observe $y_{t}$ after round $t$ is over.
The learner’s policy $\pi_{t}:[0,1]\rightarrow[0,1]$ is a mapping from points in $[0,1]$ to guesses in $[0,1]$. The optimal policy $\pi^{*}$ makes guess $\pi^{*}(\omega_{t})=1$ if $\omega_{t}>s$ and $\pi^{*}(\omega_{t})=0$ otherwise. The learner’s regret after $T$ rounds is given by
$R_{\mathrm{OLT}}(T)=T-\sum_{t=1}^{T}u_{\mathrm{OLT}}(\omega_{t},g_{t})$
since the optimal policy achieves utility $1$ in every round. In order to prove a lower bound on contextual Stackelberg regret in our setting, we make use of the following well-known lower bound on regret in the online linear thresholding setting (see e.g. [14]).
@@
Reduction from Online Linear Thresholding.
We are now ready to state our impossibility result for learning in Stackelberg games with side information whenever the sequence of contexts and the sequence of followers are chosen by an adversary.
@@
If an adversary can choose both the sequence of contexts $\mathbf{z}_{1},\ldots,\mathbf{z}_{T}$ and the sequence of followers $f_{1},\ldots,f_{T}$, no algorithm can achieve better than $\Omega(T)$ contextual Stackelberg regret in expectation over the internal randomness of the algorithm.
At a high level, the reduction from online linear thresholding proceeds by creating an instance of our setting such that the sequence of contexts $z_{1},\ldots,z_{T}$ correspond to the sequence of points $\omega_{1},\ldots,\omega_{T}$ encountered by the learner, and the sequence of follower types $f_{1},\ldots,f_{T}$ correspond to the sequence of labels $y_{1},\ldots,y_{T}$. We then show that a no-regret algorithm in the online thresholding problem can be obtained by using an algorithm which minimizes contextual Stackelberg regret on the constructed game instance as a black box. However this is a contradiction, since by Lemma 3.1 the online thresholding problem is not online learnable by any algorithm. See Figure 1 for a visualization of our reduction.

### v3

@@ Abstract
Algorithms for playing in Stackelberg games have been deployed in real-world domains including airport security, anti-poaching efforts, and cyber-crime prevention. However, these algorithms often fail to take into consideration the additional information available to each player (e.g. traffic patterns, weather conditions, network congestion), a salient feature of reality which may significantly affect both players’ optimal strategies. We formalize such settings as Stackelberg games with side information, in which both players observe an external context before playing. The leader commits to a (context-dependent) strategy, and the follower best-responds to both the leader’s strategy and the context. We focus on the online setting in which a sequence of followers arrive over time, and the context may change from round-to-round. In sharp contrast to the non-contextual version, we show that it is impossible for the leader to achieve good performance (measured by regret) in the full adversarial setting. Motivated by our impossibility result, we show that no-regret learning is possible in two natural relaxations: the setting in which the sequence of followers is chosen stochastically and the sequence of contexts is adversarial, and the setting in which the sequence of contexts is stochastic and the sequence of followers is chosen by an adversary.
@@ 1 Introduction
A Stackelberg game [31, 32] is a strategic interaction between two utility-maximizing players in which one player (the leader) is able to commit to a (possibly mixed) strategy before the other player (the follower) takes an action. While Stackelberg’s original formulation was used to model economic competition between firms, Stackelberg games have been used to study a wide range of topics in computing ranging from incentives in algorithmic decision-making [16] to radio spectrum utilization [33]. Perhaps the most successful application of Stackelberg games to solve real-world problems is in the field of security, where the analysis of Stackelberg security games has led to new methods in domains such as passenger screening at airports [8], wildlife protection efforts in conservation areas [11], the deployment of Federal Air Marshals on board commercial flights [19], and patrol boat schedules for the United States Coast Guard [1]. [^1] See [28, 20, 2] for an overview of other application domains for Stackelberg security games.
@@
We show that in the fully adversarial setting (i.e. the setting in which both the sequence of contexts and follower types is chosen by an adversary), no-regret learning is not possible, even when the policy class is highly structured. We show this via a reduction from the problem of online linear thresholding, for which it is known that no no-regret learning algorithm exists. Motivated by this impossibility result, we study two natural relaxations: (1) a setting in which the sequence of contexts is chosen by an adversary and the sequence of follower types is chosen stochastically, and (2) a setting in which the sequence of contexts is chosen stochastically and the sequence of follower types is chosen by an adversary.
@@
In the stochastic follower setting we show that the greedy algorithm (Algorithm 1), which estimates the leader’s expected utility for the given context and plays the mixed strategy which maximizes their estimate, achieves no-regret as long as the total variation distance between their estimate and the true distribution is decreasing with time. We then show how to instantiate the leader’s estimation procedure so that the regret of Algorithm 1 is $O(\min\{K,A_{f}\}\sqrt{T\log(T)})$, where $T$ is the time horizon, $K$ is the number of follower types, and $A_{f}$ is the number of follower actions. In the stochastic context setting, we show the leader can obtain $O(\sqrt{KT\log(T)}+K)$ regret by playing Hedge over a finite set of policies (Algorithm 2). An important intermediate result in both settings is that it is (nearly) without loss of generality to consider leader policies which map to a finite set of mixed strategies $\mathcal{E}_{\mathbf{z}}$, given context $\mathbf{z}$ [^2] Specifically, a leader who is restricted to playing such policies incurs negligible additional regret.
@@
Next, we extend our algorithms for both types of adversary to the setting in which the leader does not get to observe the follower’s type after each round, but instead only gets to observe their action. We refer to this type of feedback as bandit feedback. Both of our extensions to bandit feedback make use of the notion of a barycentric spanner [4], a special basis under which bounded loss estimators may be obtained for all leader mixed strategies. In the bandit stochastic follower setting, we use the fact that in addition to being bounded, a loss estimator constructed using a barycentric spanner has low variance, in order to show that a natural extension of our greedy algorithm obtains $\tilde{O}(T^{2/3})$ regret.We also make use of barycentric spanners in the (bandit) stochastic context setting, albeit in a different way. Here, our algorithm proceeds by splitting the time horizon into blocks, and using a barycentric spanner to estimate the leader’s utility from playing according to a set of special policies in each block. We then play Hedge over these policies to obtain $\tilde{O}(T^{2/3})$ leader regret. [^3] This is similar to how Balcan et al. [5] use barycentric spanners to obtain no-regret in the non-contextual setting, although more care must be taken to handle the side information present in our setting. See Table 1 for a summary of our results.
@@
Stochastic Followers, Adversarial Contexts $O\left(\min\{K,A_{f}\}\sqrt{T\log T}\right)$ (Section 4.1) $O\left(K^{2/3}A_{f}^{2/3}T^{2/3}\log^{1/3}T\right)$ (Section C.1)
Stochastic Contexts, Adversarial Followers $O\left(\sqrt{KT\log T}+K\right)$ (Section 4.2) $O\left(KA_{f}^{1/3}T^{2/3}\log^{1/3}T\right)$ (Section C.2)
@@ 1.1 Related work
Conitzer and Sandholm [9] provide algorithms for recovering the leader’s optimal mixed strategy in Stackelberg games when the follower’s payoff matrix is known to the leader. Letchford et al. [22] consider the problem of learning the leader’s optimal mixed strategy in the repeated Stackelberg game setting against a perfectly rational follower with an unknown payoff matrix. Peng et al. [24] study the same setting as Letchford et al. [22]. They provide improved rates and prove nearly-matching lower bounds. Learning algorithms to recover the leader’s optimal mixed strategy have also been studied in Stackelberg security games [5, 7, 27, 6].
Our work builds off of several results established for online learning in (non-contextual) Stackelberg games in Balcan et al. [5]. In particular, our results in Section 4.2 and Section C.2 may be viewed as a generalization of their results to the setting in which the payoffs of both players depend on an external context, in addition to the player actions. Roughly speaking, Balcan et al. [5] show that it without loss to play Hedge over a finite set of mixed strategies in order to obtain no-regret against an adversarially-chosen sequence of follower types. In order to handle the additional side information available in our setting, we instead play Hedge over a finite set of policies, each of which map to a finite set of (context-dependent) mixed strategies. However, the discretization argument is more nuanced in our setting. In particular, it is not without loss of generality to consider a finite set of policies. As a result, we need to bound the additional regret incurred by the leader due to the discretization. More recent work on learning in Stackelberg games provides improved regret rates in the full feedback [10] and bandit feedback [6] settings, and considers the effects of non-myopic followers [14] and followers who respond to calibrated predictions [15].
@@ 2 Setting and background
We consider a game between a leader and a sequence of followers. In each round $t\in[T]$, Nature reveals a context $\mathbf{z}_{t}\in\mathcal{Z}\subseteq\mathbb{R}^{d}$ to both players. [^4] E.g. in airport security, $\mathbf{z}_{t}$ may contain information about arrival and departure times, number of passengers, valuable cargo, etc. In cyber-defense, $\mathbf{z}_{t}$ may be a list of network traffic statistics. The leader moves first by playing some mixed strategy $\mathbf{x}_{t}\in\mathcal{X}\subseteq\mathbb{R}^{A}$ over a set of (finite) leader actions $\mathcal{A}$, i.e., $\mathbf{x}_{t}\in\Delta(\mathcal{A})$. The size of $\mathcal{A}$ is $A:=|\mathcal{A}|$. Having observed the leader’s mixed strategy, the follower best-responds to both $\mathbf{x}_{t}$ and $\mathbf{z}_{t}$ by playing some action $a_{f}\in\mathcal{A}_{f}$, where $\mathcal{A}_{f}$ is the (finite) set of follower actions and $A_{f}:=|\mathcal{A}_{f}|$.
@@
Follower $f$’s best-response to context $\mathbf{z}$ and mixed strategy $\mathbf{x}$ is $b_{f}(\mathbf{z},\mathbf{x})\in\arg\max_{a_{f}\in\mathcal{A}_{f}}\sum_{a_{l}\in\mathcal{A}}\mathbf{x}[a_{l}]\cdot u_{f}(\mathbf{z},a_{l},a_{f})$, where $u_{f}:\mathcal{Z}\times\mathcal{A}\times\mathcal{A}_{f}\rightarrow[0,1]$ is follower $f$’s utility function. In the case of ties, we assume that there is a fixed and known ordering over actions which determines how the follower best-responds, i.e. if $a>a^{\prime}$ for $a,a^{\prime}\in\mathcal{A}_{f}$ then the follower will break ties between $a$ and $a^{\prime}$ in favor of $a$ [^5] It is without loss of generality to assume that the follower’s best-response is a pure strategy.
We allow for the follower in round $t$ (denoted by $f_{t}$) to be one of $K\geq 1$ follower types $\{\alpha^{(1)},\ldots,\alpha^{(K)}\}$ (where $K\leq T$). Follower type $\alpha^{(i)}$ is characterized by utility function $u_{\alpha^{(i)}}:\mathcal{Z}\times\mathcal{A}\times\mathcal{A}_{f}\rightarrow[0,1]$, i.e. given a context $\mathbf{z}$, leader action $a_{l}$, and follower action $a_{f}$, a follower of type $\alpha^{(i)}$ would receive utility $u_{\alpha^{(i)}}(\mathbf{z},a_{l},a_{f})$. We assume that the set of all possible follower types and their utility functions are known to the leader, but that the follower’s type at round $t$ is not revealed to the leader until after the round is over. We denote the leader’s utility function by $u:\mathcal{Z}\times\mathcal{A}\times\mathcal{A}_{f}\rightarrow[0,1]$ and assume it is known to the leader. We often use the shorthand $u(\mathbf{z},\mathbf{x},a_{f})=\sum_{a_{l}\in\mathcal{A}}\mathbf{x}[a_{l}]\cdot u(\mathbf{z},a_{l},a_{f})$ to denote the leader’s expected utility of playing mixed strategy $\mathbf{x}$ under context $\mathbf{z}$ against follower action $a_{f}$. Follower $f_{t}$’s expected utility $u_{f_{t}}(\mathbf{z},\mathbf{x},a_{f})$ is defined analogously.
A leader policy $\pi:\mathcal{Z}\rightarrow\mathcal{X}$ is a (possibly random) mapping from contexts to mixed strategies. If the leader using policy $\pi_{t}$ in round $t$ and observes context $\mathbf{z}_{t}$, their strategy $\mathbf{x}_{t}$ is given by $\mathbf{x}_{t}\sim\pi_{t}(\mathbf{z}_{t})$.
@@
If an algorithm achieves regret $R(T)=o(T)$, we say that it is a no-regret algorithm. We consider three ways in which Nature can select the sequence of contexts/followers:
@@
    If Nature chooses context $\mathbf{z}_{t}$ (resp. follower $f_{t}$) before round $t$ in order to harm the leader (possibly using knowledge of the leader’s algorithm and the outcomes of the prior $t-1$ rounds), we say that the sequence of contexts (resp. follower types) are chosen by an adaptive adversary.
@@ 3 On the impossibility of fully adversarial no-regret learning
We begin with a negative result: no-regret learning is not possible in the setting of Section 2 if the sequence of contexts and the sequence of followers is chosen by an adversary. While this is not necessarily surprising given that Definition 2.3 allows for the optimal policy $\pi^{*}$ to be arbitrarily complex, we show that this result holds even when the policy class to which $\pi^{*}$ belongs is highly structured. We show this via a reduction to the online linear thresholding problem, for which it is known that no-regret learning is impossible.
Online linear thresholding.
The online linear thresholding problem is a repeated two-player game between a learner and an adversary. Before the first round, an adversary chooses a cutoff $s\in[0,1]$ which is unknown to the learner. In each round, the adversary chooses a point $\omega_{t}\in[0,1]$ and reveals it to the learner. $\omega_{t}$ is assigned label $y_{t}=1$ if $\omega_{t}>s$ and label $y_{t}=-1$ otherwise. Given $\omega_{t}$, the learner makes a guess $g_{t}\in[0,1]$ (the probability they place on $y_{t}=1$), and receives utility $u_{\mathrm{OLT}}(\omega_{t},g_{t})=g_{t}\cdot\mathbbm{1}\{y_{t}=1\}+(1-g_{t})\cdot\mathbbm{1}\{y_{t}=-1\}$. The learner gets to observe $y_{t}$ after round $t$ is over. The learner’s policy $\pi_{t}:[0,1]\rightarrow[0,1]$ is a mapping from points in $[0,1]$ to guesses in $[0,1]$. The optimal policy $\pi^{*}$ makes guess $\pi^{*}(\omega_{t})=1$ if $\omega_{t}>s$ and $\pi^{*}(\omega_{t})=0$ otherwise. The learner’s regret after $T$ rounds is given by $R_{\mathrm{OLT}}(T)=T-\sum_{t=1}^{T}u_{\mathrm{OLT}}(\omega_{t},g_{t})$, since the optimal policy achieves utility $1$ in every round. In order to prove a lower bound on contextual Stackelberg regret in our setting, we make use of the following well-known lower bound on regret in the online linear thresholding setting (see e.g. [13]).
@@
If an adversary can choose both the sequence of contexts $\mathbf{z}_{1},\ldots,\mathbf{z}_{T}$ and the sequence of followers $f_{1},\ldots,f_{T}$, no algorithm can achieve better than $\Omega(T)$ contextual Stackelberg regret in expectation over the internal randomness of the algorithm, even when $\pi^{*}$ is restricted to come from the set of linear thresholding functions.
The reduction from online linear thresholding proceeds by creating an instance of our setting such that the sequence of contexts $z_{1},\ldots,z_{T}$ correspond to the sequence of points $\omega_{1},\ldots,\omega_{T}$ encountered by the learner, and the sequence of follower types $f_{1},\ldots,f_{T}$ correspond to the sequence of labels $y_{1},\ldots,y_{T}$. We then show that a no-regret algorithm in the online thresholding problem can be obtained by using an algorithm which minimizes contextual Stackelberg regret on the constructed game instance as a black box. However this is a contradiction, since by Lemma 3.1 the online thresholding problem is not online learnable by any algorithm. See Figure 1 for a visualization of our reduction.